OpenStudy (anonymous):

3^(1-2x) = 4^x Solve with logs.

OpenStudy (anonymous):

\[3^{1-2x} = 4^{x}\]

OpenStudy (anonymous):

Take the natural log of both sides to get rid of the exponents. This turns the exponents into factors. (1-2x)ln(3) = xln(4) Then, solve for x.